OpenStudy (anonymous):

2m²+27m+70 a.(2m+7)(m+10) b.(2m+10)(m+7) c.(2m+3)(m+20) d.(2m+7)(2m-10)

OpenStudy (anonymous):

How do you factor this?

Directrix (directrix):

2m²+27m+70 What are two numbers that multiply to 2*70 and also add to 27 ?

OpenStudy (anonymous):

i dont think any of the answers do. Thats whats messing me up

Directrix (directrix):

Did you get the factors of 140?

Directrix (directrix):

Check out this online prime factor finder. Just enter 140. http://www.calculatorsoup.com/calculators/math/prime-factors.php

Directrix (directrix):

How about 7 and 20? They multiply to 140 and add to 27.

OpenStudy (anonymous):

but i thought we needed them to multiply to 70

Directrix (directrix):

On the first problem we did together, the coefficient of the x^2 term was 1. And, 1 times whatever the last number was yielded just the last number. -------- The numbers we want are the numbers that multiply to a*c and add to b. In this problem, a = 2 and c = 70. b = 27. We want two numbers that multiply to 140 and add to 27. Just remember: Multiply to ac and add to b.

Directrix (directrix):

We cannot quit now.

Directrix (directrix):

By the way, it took me two weeks to learn to factor.

Directrix (directrix):

2m²+27m+70 = 2m^2 + 20m + 7m + 70

Directrix (directrix):

2m^2 + 20m + 7m + 70 = 2m ( m + 10) + 7(m + 10)

Directrix (directrix):

2m ( m + 10) + 7(m + 10) = (m + 10) (2m + 7)

Directrix (directrix):

@JaneDoe100

OpenStudy (anonymous):

@Directrix sorry im watching my little sister and had to put her down for a nap. i think i understand it now

Directrix (directrix):

@JaneDoe100 No worries. Don't give up on learning to factor.

OpenStudy (anonymous):

ok i wont